OpenStudy (wesdg1978):

For the initial value problem: dy/dx=4e^x, y(0)=1 I did integral 4e^x dx which equals 4e^x+c , for y(0)=1 , what do I do next? The answer is y=4e^x-3

OpenStudy (zzr0ck3r):

solve for c y = 4e^x+c you have the point (0,1) 1 = 4e^0+c c = -3 so y=4e^x - 3

OpenStudy (wesdg1978):

I don't see what I'm missing, I don't understand how c=-3, could you explain that step a little more please?

OpenStudy (anonymous):

$$y(x)=4e^x+c$$

OpenStudy (anonymous):

$$y(0)=4+c$$

OpenStudy (anonymous):

$$1=4+c$$

OpenStudy (anonymous):

$$c=-3$$

OpenStudy (zzr0ck3r):

ok you have the point (0,1) that is when x = 0 you have y = 1 so let y = 1, and x = 0 1 = 4e^(0) + c 1=4+c c=-3

OpenStudy (wesdg1978):

so 4e^0 isn't 4(0) which is 0 ?

OpenStudy (zzr0ck3r):

its like if I gave you the line y = 3x+b and told you that the point (0,3) lies on the line

OpenStudy (zzr0ck3r):

correct, x = 0

OpenStudy (wesdg1978):

Thanks!